Find the diameter of copper wire with which it has the same resistance as aluminium wire of equal length and diameter 1.20mm. The resistivity of copper and aluminum at room temperature are 1.7×10^-8ohms per meter and 2.6× 10^-8ohms per meter respectively


1
Expert's answer
2021-03-31T09:53:12-0400



To be given in question

RAl=Rcu

LAl=Lcu

Aluminium wire diameter D(Al)=1.20mm

Resistivity copper wire"\rho" (cu)=1.7"\times" "10^{-8}" ohm/meter

Resistivity aluminium wire "\rho" (Al)=2.6"\times" "10^{-8}" ohm/meter

To be asked in question


Diameter of copper wire D(cu)=?

Formula "R=\frac{\rho L}{A}"


=RAl=Rcu

["\frac{\rho L}{A}" ]Al=[ "\frac{\rho L}{A}" ] cu

Both wire length are equal

"[\frac{\rho }{A}] (Al) =[\frac{\rho }{A}](cu)"

Put value 

"[\frac{2.6\times10^{-8}\times4}{\pi\times[ 1.20\times 10^{-3}]^2}]= [\frac{1.7\times10^{-8}\times4}{\pi\times [D^{2}(cu)]}]"

"= 9.703 \times 10^{-04} meter"